OpenStudy (anonymous):

Use rational exponents to write, cbrt(x)*^5√(9x) as a single radical expression.

OpenStudy (amistre64):

convert the index to exponentials

OpenStudy (amistre64):

3rt = ^1/3 4rt = ^1/4 nrt = ^1/n

OpenStudy (amistre64):

that way youve got it in exponentials of xs that are easier to see and work with

OpenStudy (amistre64):

\[\sqrt[3]{x}*\sqrt[5]{9x}\] \[ x^{1/3} *9^{1/5}x^{1/5}\] \[ 9^{1/5}x^{1/5+1/3}\] etc ...